OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^2 + 4x + 7

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{4}{2}=-2\] so start with \[y=(x+2)^2+k\] and to find \(k\) replace \(x\) by \(-2\) and see what you get

OpenStudy (anonymous):

@satellite73 I got: y = 0 + k I'm sort of confused, because here are my answer choices: y = (x + 2)^2 + 3 y = (x + 2)^2 - 3 y = (x - 2)^2 - 3 y = (x - 2)^2 + 3

OpenStudy (anonymous):

@satellite73 Can you help me please??? (: